OpenStudy (anonymous):

If two dies are toced together what is the probability of having sum of 5 as the outcome

OpenStudy (anonymous):

favorable ways:(1,4),(4,1),(2,3),(3,2) four ways total outcomes are 6*6=36 divide favorable by total and get result

OpenStudy (anonymous):

4/36=1/9